The sum of Mary's age and twice Ann's age is 43. Twice Mary's age is 11 more than Ann's age. How old are they?
Answer by stanbon(75887)

The sum of Mary's age and twice Ann's age is 43. Twice Mary's age is 11 more than Ann's age. How old are they?
-----------------
M + 2A = 43
2M = A + 11
----------------
Substitute for "A" and solve for "M"::
M + 2(2M-11) = 43
M + 4M - 22 = 43
5M = 65
Mary = 13 yrs
----
A = 2M-11 = 26-11 = 15 yrs (Ann's age)
-------------
Cheers,
Stan H.
-------------